OpenStudy (anonymous):

Find all the zeros of each polynomial function. f(x) = x4 + 4x3 – 6x2 – 36x – 27 Select one: a. ±1 and ±3 b. ±3 only c. ±1 only d. none of the above

OpenStudy (anonymous):

@TylerD

OpenStudy (zpupster):

try the factors of the coefficient + and - and substitute them into the eqn. those with an answer of 0 are the zeros 1.f(x)=x^4+4x^3-6x^2-36x-27 f(1)=1^4+4(1)^3-6(1)^2-36(1)-27 = -64 not a zero f(-1)= (-1)^4+4(-1)^3-6(-1)^2-36(-1)-27 = 0 thus zero x = -1 is a zero so (x+1) is a factor of the polynomial if you look at your answers 1 is not a zero s0 ??

OpenStudy (zpupster):

but -1 is

OpenStudy (anonymous):

so which answer are we looking at

OpenStudy (anonymous):

@TylerD

OpenStudy (anonymous):

since you are given options, perhaps the easiest thing to do would be to check

OpenStudy (anonymous):

\[f(x) = x^4 + 4x^3 – 6x^2 – 36x – 27 \] \[f(1) = 1 + 4– 6 – 36 – 27 \neq 0\] so that eliminates A and C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it D non of the above